The sum of ages of a woman and her daughter is 46 years in 4 years time, the ratios of their age will be 7:2, find the present age
Answer by umnikmo(75)

X-present age of the woman
Y-present age of the daughter
X+Y=46 => X=46-Y (1)
(X+4)/(Y+4)=7/2 => 2(X+4)=7(Y+4) => 2X+8=7Y+28 => X=(7Y+20)/2 (2)
NOW (1)=(2):
46-Y=(7Y+20)/2
92-2Y=7Y+20
72=9Y
Y=8
X=46-8=38